Paloma Rey drove to work on Wednesday at 40 miles per hour, an darrived one minute late. She left home at the same time on Thursday, drove 45 miles per hour, and arrived one minute early. How far does Ms. Rey drive to work?
Answer by jayanthi(5)

Let time be X hrs
Speed * Time = Distance
= 40(X+1/60) = 45(X-1/60)
= 40X = 2/3 = 45X - 3/4
= 5X = 2/3 + 3/4
= 8+9/12 = 17/12
= X = 17/60 hrs = 17mints
= Distance = Speed * Time
= 40(17/60+1/60)
= 40 * 18/60

= 12miles